# NEP: Autoregressive Image Editing via Next Editing Token Prediction

### v1

@@ Abstract
Text-guided image editing involves modifying a source image based on a language instruction and, typically, requires changes to only small local regions. However, existing approaches generate the entire target image rather than selectively regenerate only the intended editing areas. This results in (1) unnecessary computational costs and (2) a bias toward reconstructing non-editing regions, which compromises the quality of the intended edits. To resolve these limitations, we propose to formulate image editing as Next Editing-token Prediction (NEP) based on autoregressive image generation, where only regions that need to be edited are regenerated, thus avoiding unintended modification to the non-editing areas. To enable any-region editing, we propose to pre-train an any-order autoregressive text-to-image (T2I) model. Once trained, it is capable of zero-shot image editing and can be easily adapted to NEP for image editing, which achieves a new state-of-the-art on widely used image editing benchmarks. Moreover, our model naturally supports test-time scaling (TTS) through iteratively refining its generation in a zero-shot manner. The project page is: https://nep-bigai.github.io/
Figure 1: Our approach avoids full-image generation and does not introduce unintended changes as the previous diffusion-model-based editing approach Zhang et al. (2024).
@@ 1 Introduction
Text-driven image editing aims to modify a source image following a given language instruction. Typically, modifications are confined to small local regions (editing regions) while most of the image remains unchanged (non-editing regions). A predominant paradigm for solving the task is through the diffusion model Song et al. (2022, 2021); Rombach et al. (2022a), but standard diffusion models struggle with controllable editing, that is, editing only a target region without altering the surrounding areas. To tackle this challenge, an inversion technique has been proposed and augmented with diffusion-based image generation models Song et al. (2021); Mokady et al. (2023). The core idea of this method is to find the mapping of non-editing regions to the corresponding subspace of Gaussian noise. It requires that the initial Gaussian noise that can be decoded into the source image should be pre-defined, which is, however, hard to obtain exactly, and further leads to unintended edits Hertz et al. (2022).
A more controllable paradigm is to pre-define editing regions and edit only the specified areas while preserving the rest Avrahami et al. (2022); Nichol et al. (2022). However, these approaches perform full generation of the target image, including regions that are not required to be edited, and thus are suboptimal in terms of efficiency. This inefficiency is pronounced in training-based editing approaches Brooks et al. (2023); Xiao et al. (2024), which also demand significant computational resources to learn to reconstruct. Moreover, the reliance on full-image generation introduces a learning bias during training; that is, image editing models tend to prioritize reconstruction for the non-editing regions over regeneration for the intended editing regions Xiao et al. (2024).
@@ 2.1 NEP Pre-training
Preliminaries on LlamaGen.   The NEP framework is versatile and compatible with various design choicesSun et al. (2024); Wu et al. (2024); Wang et al. (2024). In this work, we build upon LlamaGenSun et al. (2024), the first open-source text-conditioned autoregressive model to outperform diffusion models, leveraging its robust architecture to enable NEP’s random-order generation and iterative refinement for enhanced image editing and generation. To maintain potential unification with text modality, the architecture design of LlamaGen largely follows one of the popular LLMs, Llama Touvron et al. (2023a, b). The conditioning text embeddings are extracted from FLAN T5 Chung et al. (2024), followed by a projector for dimensionality alignment. The text embeddings are left-padded to a fixed length $L_{T}$ and prefilled to generate image tokens. Images are firstly tokenized by the encoder and quantizer of VQGAN Esser et al. (2021), and generated token ids are mapped to RGB pixels by the decoder. Image tokens with length $L$ are generated in a next-token prediction fashion. Formally, given a text sequence $T$, the sequentialized image tokens $I=\{I_{1},I_{2},...,I_{L}\}$, are generated by:
@@
Figure 2: Overview of Next-Editing-token Prediction. The input sequence is comprised of: 1) text embeddings, extracted from FLAN-T5, 2) source image embeddings, tokenized by VQGAN, and 3) mask embeddings, a sequence of interleaved editing and non-editing embeddings. The output editing tokens (in raster scan order) are filled back to the source image based on the editing mask.
RLlamaGen: Randomized Autoregressive Text-to-Image Generation.   To address LlamaGen’s limitation of generating image tokens solely in raster scan order, we extend it to create RllamaGen, which supports generating image tokens in any user-specified order, enabling flexible, arbitrary-order generation. To add order awareness to the model, following Pannatier et al. (2024); Yu et al. (2024), we learn an extra sequence of positional embeddings $PE_{1},PE_{2},...,PE_{L}$, which is shuffled based on a random order to define the generation sequence. For each input image token, the positional embedding corresponding to the next token in the assigned order is added. Formally, the generation of an image sequence $I_{O}$ in the order of $O=[o_{1},o_{2},...o_{L}]$ is defined as:
@@ 2.2 NEP: Next-Editing-token Prediction
Editing Region Conditioning We firstly patchify the pixel-level editing mask $M$ by max-pooling each non-overlapping sliding window with the size of $p\times p$. Subsequently, we flatten the patched mask into a sequence $M^{E}=\{m_{1},m_{2},...m_{L}\}\in\{0,1\}^{L}$. The masking sequence $I^{M}=\{I^{M}_{1},...,I^{M}_{L}\}$ is tokenized by querying a two-sized codebook comprising an editing embedding $E_{emb}$ and a non-editing embedding $U_{emb}$, which is formally defined as:
@@ 2.3 Test-time Scaling with NEP
NEP can be employed to support test-time scaling by integrating it into a self-improving loop. In each refinement step, prior to NEP, a revision region is proposed. Existing image reward models  Xu et al. (2023) usually produce a single value for the full image. To obtain token-level dense quality scores, we calculate Grad-CAM Selvaraju et al. (2017) value regarding the critic model. These values reflect each token’s contribution to the overall image quality score, measured by a reward model. Positions that correspond to the $K$ lowest scores are identified as the revision regions. During revision, we adopt NEP to regenerate tokens in this region, conditioning them on the remaining high-quality tokens. After NEP, a reward model evaluates whether the revised image surpasses the original, determining whether to accept or reject the revision. To further improve quality, for NEP, we apply a rejection sampling strategy, regenerating tokens at the revision positions in multiple random orders and selecting the revision with the highest quality score. This approach demonstrates strong scaling potential, suggesting that effective revision of initial generations can significantly enhance performance.
@@ 3.1 Datasets and Training settings
T2I pre-training settings.   We use LlamaGen-XL with 775M parameters as the base T2I model and adapt it to RLlamaGen adding 0.3M positional embedding parameters. Our training data consists of around 16M text-image pairs and is collected from multiple open-source datasets, including ALLaVA-LAION Chen et al. (2024a), CC12M Changpinyo et al. (2021), Kosmos-G Pan et al. (2024), LAION-LVIS-220 Schuhmann and Bevan (2023), LAION-COCO-AESTHETIC Liu (2023), LAION-COCO-17M Zhang (2024), and ShareGPT4V Chen et al. (2024b). We train RLlamaGen for $60,000$ steps with a batch size of $360$ and an image resolution of $256\times 256$. The optimizer is Fused AdamW with $\beta_{1}$, $\beta_{2}$ set to 0.9, 0.95, respectively, and a constant learning rate of 1e-4 is used. We perform training on 8 NVIDIA Tesla A100 GPUs, which takes 39 hours.
Image Editing Training Settings.   We fine-tune RLlamaGen for image editing by adding two learnable embeddings (i.e., $E_{emb}$ and $U_{emb}$) to specify masking regions. This strategy is computationally efficient, with only 3.6k parameters introduced. Our editing model is trained on the UltraEdit dataset Zhao et al. (2024) that comprises 4 million image pairs, where 131k samples are annotated with editing regions. For those with no editing region annotations, we use them for full-image generation.
We perform training on 4 NVIDIA Tesla A100 GPUs. The model is trained for $3.9M$ steps with a batch size of $100$ and a learning rate of 1e41e-4. Per common practices Zhang et al. (2024); Zhao et al. (2024), we evaluate models at a higher image resolution than that used during training (specifically, $512\times 512$ pixels compared to $256\times 256$ pixels), and fine-tune them on the target resolution for an additional $2,000$ steps. For the Emu Edit benchmark, we train our model with a learning rate of 1e51e-5 for $60,000$ steps.
@@ 3.2 Results on Image Editing
Benchmarks & Evaluation Metrics.   We demonstrate the superiority of our approach on two widely recognized benchmarks: MagicBrush Zhang et al. (2024) and Emu Edit Sheynin et al. (2023). The MagicBrush test set provides editing region annotations for each sample, thereby facilitating the evaluation of region-conditioned editing. This benchmark assesses both multi-turn editing, which evaluates the final image after a series of edits, and single-turn editing, which assesses the target image following an individual edit.
@@
3.2.1 Quantitative Results
We demonstrate the superiority of NEP in terms of region-aware editing on the MagicBrush test set. The compared prior arts broadly fall into two categories: (1) global description-based, such as SD-SDEdit (Meng et al., 2022), Null Text Inversion (Mokady et al., 2023), GLIDE (Nichol et al., 2022), as well as Blended Diffusion (Avrahami et al., 2022), and (2) instruction-guided, including HIVE (Zhang et al., 2023), InstructPix2Pix (Brooks et al., 2023), MagicBrush (Zhang et al., 2024), UltraEdit (Zhao et al., 2024), FireEdit Zhou et al. (2025), AnySD Yu et al. (2025) and EditAR Mu et al. (2025). Table 1 demonstrates that our approach achieves the highest score for single-turn editing and better or comparable performance under the multi-turn setting. For the first time, autoregressive models can achieve top performance on well-recognized editing benchmarks.
We demonstrate the effectiveness of free-form editing on the Emu Edit test set Sheynin et al. (2023). We compare NEP with state-of-the-art approaches including InstructPix2Pix (Brooks et al., 2023), MagicBrush (Zhang et al., 2024), Emu Edit Sheynin et al. (2023) UltraEdit (Zhao et al., 2024), MIGE Tian et al. (2025), and AnySD Yu et al. (2025). In Table 2, we can observe that, without resorting to editing masks, our approach still achieves comparable or better editing performance.
3.2.2 Ablation Study
We perform ablation studies on the Magicbrush multi-turn test set. For each configuration, we report the results of the models trained for $30,000$ steps. We assess two critical design choices. First, we exclude mask embeddings, relying solely on text and source images as inputs, which degrades performance as shown in Table 3. Second, we remove the next editing token positions by generating all tokens in a raster scan order, following an NTP framework. Without any priors on editing regions, this leads to a significant performance drop, highlighting the need for targeted token generation.
Table 3: Ablation studies on the MagicBrush test set under the multi-turn setting. We validate the contribution of each design choice by removing them and observing the performance drop. We ablate two aspects: 1) Editing region conditioning by removing the editing & unediting tokens inferred from editing region masks, and 2) NEP vs. NTP by generating full image tokens. The default setting is highlighted in gray.
@@ 3.3.1 Zero-shot Image Editing
We demonstrate that RLlmaGen is readily capable of image editing. This is achieved by removing the tokens in the editing region and regenerating them. Figure 4 demonstrates that RLlmaGen can make fine-grained and coherent edits.
@@
3.3.2 Text-to-Image Generation Results
Benchmarks & Evaluation Metrics.   We evaluate the image generation quality on MS-COCO 30K in terms of Fréchet Inception Distance (FID) and CLIP similarity. FID reflects the fidelity and diversity of generated images. It measures the distance between the ground truth image distribution and the generated image distribution, where the distributions are constituted of Inception V3 Szegedy et al. (2016) embeddings extracted from corresponding images. The CLIP score is used to evaluate the instruction-following ability of T2I models. It measures the similarity between the vision embeddings extracted from the generated image and text encoder embeddings extracted from corresponding captions.
@@ 4.1 Text-to-Image Generation
Text-to-image generation has become a cornerstone of modern artificial intelligence, enabling to create visual content based on textual descriptions. Pioneering models such as Generative Adversarial Networks (GANs) Goodfellow et al. (2020) make groundbreaking breakthroughs by generating high-fidelity images. AttnGAN Xu et al. (2018) built on StackGAN Zhang et al. (2017) achieves better alignment with text instructions. However, GANs still faced challenges like training instability (e.g., mode collapse, where the model generates limited varieties of images) and difficulty with highly detailed or multi-object scenes, setting the stage for the next evolutionary step.
More recently, diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Song et al. (2020a) like Stable Diffusion Rombach et al. (2022b) have emerged, creating realistic images by iteratively denoising random noise guided by text descriptions, setting a new standard for quality and versatility. However, the learning paradim and architectures diverge from well-established large language models (LLMs) Brown et al. (2020),, making it difficult for artificial general intelligence featuring a shared framework for various modalities.
In this regard, a line of works Ramesh et al. (2021, 2022); Yu et al. (2022) resort to autoregressive models for visual generation. Images are tokenized into a sequence of tokens and generated sequentially based on prefilled text tokens. Benefiting from large-scale models and datasets, they can create photorealistic images with a remarkable text-following capability. This field is further advanced by several open-source works, such as LlamaGen Sun et al. (2024), Emu3 Wang et al. (2024), and Janus Wu et al. (2024).

### v2

@@ Abstract
Text-guided image editing involves modifying a source image based on a language instruction and, typically, requires changes to only small local regions. However, existing approaches generate the entire target image rather than selectively regenerate only the intended editing areas. This results in (1) unnecessary computational costs and (2) a bias toward reconstructing non-editing regions, which compromises the quality of the intended edits. To resolve these limitations, we propose to formulate image editing as Next Editing-token Prediction (NEP) based on autoregressive image generation, where only regions that need to be edited are regenerated, thus avoiding unintended modification to the non-editing areas. To enable any-region editing, we propose to pre-train an any-order autoregressive text-to-image (T2I) model. Once trained, it is capable of zero-shot image editing and can be easily adapted to NEP for image editing, which achieves a new state-of-the-art on widely used image editing benchmarks. Moreover, our model naturally supports test-time scaling (TTS) through iteratively refining its generation in a zero-shot manner.
Figure 1: Our approach avoids full-image generation and does not introduce unintended changes as the previous diffusion-model-based editing approach [58].
@@ 1 Introduction
Text-driven image editing aims to modify a source image following a given language instruction. Typically, modifications are confined to small local regions (editing regions) while most of the image remains unchanged (non-editing regions). A predominant paradigm for solving the task is through the diffusion model [40, 39, 31], but standard diffusion models struggle with controllable editing, that is, editing only a target region without altering the surrounding areas. To tackle this challenge, an inversion technique has been proposed and augmented with diffusion-based image generation models [39, 21]. The core idea of this method is to find the mapping of non-editing regions to the corresponding subspace of Gaussian noise. It requires that the initial Gaussian noise that can be decoded into the source image should be pre-defined, which is, however, hard to obtain exactly, and further leads to unintended edits [11].
A more controllable paradigm is to pre-define editing regions and edit only the specified areas while preserving the rest [1, 23]. However, these approaches perform full generation of the target image, including regions that are not required to be edited, and thus are suboptimal in terms of efficiency. This inefficiency is pronounced in training-based editing approaches [2, 50], which also demand significant computational resources to learn to reconstruct. Moreover, the reliance on full-image generation introduces a learning bias during training; that is, image editing models tend to prioritize reconstruction for the non-editing regions over regeneration for the intended editing regions [50].
@@ 2.1 NEP Pre-training
Preliminaries on LlamaGen.   The NEP framework is versatile and compatible with various design choices[42, 49, 48]. In this work, we build upon LlamaGen[42], the first open-source text-conditioned autoregressive model to outperform diffusion models, leveraging its robust architecture to enable NEP’s random-order generation and iterative refinement for enhanced image editing and generation. To maintain potential unification with text modality, the architecture design of LlamaGen largely follows one of the popular LLMs, Llama [45, 46]. The conditioning text embeddings are extracted from FLAN T5 [7], followed by a projector for dimensionality alignment. The text embeddings are left-padded to a fixed length $L_{T}$ and prefilled to generate image tokens. Images are firstly tokenized by the encoder and quantizer of VQGAN [8], and generated token ids are mapped to RGB pixels by the decoder. Image tokens with length $L$ are generated in a next-token prediction fashion. Formally, given a text sequence $T$, the sequentialized image tokens $I=\{I_{1},I_{2},...,I_{L}\}$, are generated by:
@@
Figure 2: Overview of Next-Editing-token Prediction. The input sequence is comprised of: 1) text embeddings, extracted from FLAN-T5, 2) source image embeddings, tokenized by VQGAN, and 3) mask embeddings, a sequence of interleaved editing and non-editing embeddings. The output editing tokens (in raster scan order) are filled back to the source image based on the editing mask. $PE_{i}$ denotes the learned positional embeddings that specify the token generation order.
RLlamaGen: Randomized Autoregressive Text-to-Image Generation.   To address LlamaGen’s limitation of generating image tokens solely in raster scan order, we extend it to create RllamaGen, which supports generating image tokens in any user-specified order, enabling flexible, arbitrary-order generation [26, 55, 25]. To add order awareness to the model, following [26, 55], we learn an extra sequence of positional embeddings $PE_{1},PE_{2},...,PE_{L}$, which is shuffled based on a random order to define the generation sequence. For each input image token, the positional embedding corresponding to the next token in the assigned order is added. Formally, the generation of an image sequence $I_{O}$ in the order of $O=[o_{1},o_{2},...o_{L}]$ is defined as:
@@ 2.2 NEP: Next-Editing-token Prediction
Editing Region Conditioning (ERC) We firstly patchify the pixel-level editing mask $M$ by max-pooling each non-overlapping sliding window with the size of $p\times p$. Subsequently, we flatten the patched mask into a sequence $M^{E}=\{m_{1},m_{2},...m_{L}\}\in\{0,1\}^{L}$. The masking sequence $I^{M}=\{I^{M}_{1},...,I^{M}_{L}\}$ is tokenized by querying a two-sized codebook comprising an editing embedding $E_{emb}$ and a non-editing embedding $U_{emb}$, which is formally defined as:
@@ 2.3 Test-time Scaling with NEP
NEP can be employed to support test-time scaling by integrating it into a self-improving loop. In each refinement step, prior to NEP, a revision region is proposed. Existing image reward models  [51] usually produce a single value for the full image. To obtain token-level dense quality scores, we calculate Grad-CAM [35] value regarding the critic model (i.e., off-the-shelf CLIP-ViT-B/32). These values reflect each token’s contribution to the overall image quality score, measured by a reward model (i.e., ImageReward [51]). Positions that correspond to the $K$ lowest scores are identified as the revision regions. During revision, we adopt NEP to regenerate tokens in this region, conditioning them on the remaining high-quality tokens. After NEP, the reward model evaluates whether the revised image surpasses the original, determining whether to accept or reject the revision. To further improve quality, for NEP, we apply a rejection sampling strategy, regenerating tokens at the revision positions in multiple random orders and selecting the revision with the highest quality score. This approach demonstrates strong scaling potential, suggesting that effective revision of initial generations can significantly enhance performance.
@@ 3.1 Datasets and Training settings
T2I pre-training settings.   We use LlamaGen-XL with 775M parameters as the base T2I model and adapt it to RLlamaGen adding 0.3M positional embedding parameters. Our training data consists of around 16M text-image pairs and is collected from multiple open-source datasets, including ALLaVA-LAION [5], CC12M [4], Kosmos-G [24], LAION-LVIS-220 [34], LAION-COCO-AESTHETIC [18], LAION-COCO-17M [56], and ShareGPT4V [6]. We train RLlamaGen for $60,000$ steps with a batch size of $360$ and an image resolution of $256\times 256$. The optimizer is Fused AdamW with $\beta_{1}$, $\beta_{2}$ set to 0.9, 0.95, respectively, and a constant learning rate of 1e-4 is used. We perform training on 8 NVIDIA Tesla A100 GPUs, which takes 39 hours.
Image Editing Training Settings.   We fine-tune RLlamaGen for image editing by adding two learnable embeddings (i.e., $E_{emb}$ and $U_{emb}$) to specify masking regions. This strategy is computationally efficient, with only 3.6k parameters introduced. Our editing model is trained on the UltraEdit dataset [60] that comprises 4 million image pairs, where 131k samples are annotated with editing regions. For those with no editing region annotations, we use them for full-image generation.
We perform training on 4 NVIDIA Tesla A100 GPUs. The model is trained for $3.9M$ steps with a batch size of $100$ and a learning rate of 1e41e-4. Per common practices [58, 60], we evaluate models at a higher image resolution than that used during training (specifically, $512\times 512$ pixels compared to $256\times 256$ pixels), and fine-tune them on the target resolution for an additional $2,000$ steps. For the Emu Edit benchmark, we train our model with a learning rate of 1e51e-5 for $60,000$ steps.
@@ 3.2 Results on Image Editing
Benchmarks & Evaluation Metrics.   We demonstrate the superiority of our approach on two widely recognized benchmarks: MagicBrush [58] and Emu Edit [36]. The MagicBrush test set provides editing region annotations for each sample, thereby facilitating the evaluation of region-conditioned editing. This benchmark assesses both multi-turn editing, which evaluates the final image after a series of edits, and single-turn editing, which assesses the target image following an individual edit.
@@
3.2.1 Quantitative Results
We demonstrate the superiority of NEP in terms of region-aware editing on the MagicBrush test set. The compared prior arts broadly fall into two categories: (1) global description-based, such as SD-SDEdit [20], Null Text Inversion [21], GLIDE [23], as well as Blended Diffusion [1], and (2) instruction-guided, including HIVE [59], InstructPix2Pix [2], MagicBrush [58], UltraEdit [60], FireEdit [61], AnySD [54] and EditAR [22]. Table 1 demonstrates that our approach achieves the highest score for single-turn editing and better or comparable performance under the multi-turn setting. For the first time, autoregressive models can achieve top performance on well-recognized editing benchmarks.
@@ 3.2.1 Quantitative Results
Table 3: Ablation studies on the MagicBrush test set under the multi-turn setting. We validate the contribution of each design choice by removing them and observing the performance drop. We ablate two aspects: 1) ERC by removing the editing & unediting tokens inferred from editing region masks, and 2) NEP vs. NTP by generating full image tokens. The default setting is highlighted in gray.
@@
We demonstrate the effectiveness of free-form editing on the Emu Edit test set [36]. We compare NEP with state-of-the-art approaches including InstructPix2Pix [2], MagicBrush [58], Emu Edit [36] UltraEdit [60], MIGE [44], and AnySD [54]. In Table 2, we can observe that, without resorting to editing masks, our approach still achieves comparable or better editing performance.
Figure 3: Visualized ablation on ERC.  This demonstrates that removing Editing Region Conditioning increases the editing model’s change to refuse to modify the source image. Best viewed zoomed in and in color.
3.2.2 Ablation Study
We perform ablation studies on the Magicbrush multi-turn test set. For each configuration, we report the results of the models trained for $30,000$ steps. We assess two critical design choices. First, we exclude mask embeddings, relying solely on text and source images as inputs, which degrades performance as shown in Table 3. Qualitatively, we observe that removing ERC increases the likelihood of the model making no changes to the source model, as demonstrated in Figure 3. Second, we remove the next editing token positions by generating all tokens in a raster scan order, following an NTP framework. Without any priors on editing regions, this leads to a significant performance drop, highlighting the need for targeted token generation.
@@ 3.3.1 Zero-shot Image Editing
We demonstrate that RLlmaGen is readily capable of image editing. This is achieved by regenerating tokens in the editing regions. Figure 5 demonstrates that RLlmaGen can make fine-grained and coherent edits.
Table 5: Comparative Results on Zero-shot Editing on MagicBrush test set.
[…]
Single-turn aMUSEd 0.0913 0.0300 0.8802 0.8131
Ours (zero-shot) 0.0743 0.0211 0.9032 0.8509
Multi-turn aMUSEd 0.1034 0.0361 0.8689 0.8092
Ours (zero-shot) 0.0916 0.0319 0.8798 0.7859
Comparison with Localized Editing Approaches.  We compare our zero-shot editing performance against aMUSEd [27], which is also capable of localized zero-shot editing. We use its publicly available checkpoint for comparison, adhering to its default configurations [^1] https://huggingface.co/blog/amused. Results on the MagicBrush dataset show that our approach outperforms aMUSEd. This is attributed to our method’s ability to enable fine-grained editing by keeping all source image tokens visible to the generation model, whereas aMUSEd replaces edited regions with mask tokens, limiting its precision.
Table 6: Ablations on Generation Order for Zero-shot Editing on MagicBrush test set.
[…]
Single-turn In-mask random order 0.0741 0.0211 0.9027 0.8482
In-mask raster scan order 0.0743 0.0211 0.9032 0.8509
Multi-turn In-mask random order 0.0911 0.0316 0.8782 0.7833
In-mask raster scan order 0.0916 0.0319 0.8798 0.7859
Ablations on Generation Order.  Alternative to the default generation order, i.e., an in-mask raster scan order, as we introduced in Section 3.3, we employ random generation order for zero-shot image editing. The results in Table 6 demonstrate that altering the generation order has negligible impact on the effectiveness of our approach, confirming its robustness.
3.3.2 Text-to-Image Generation Results
Benchmarks & Evaluation Metrics.   We evaluate the image generation quality on MS-COCO 30K in terms of Fréchet Inception Distance (FID) and CLIP similarity. FID reflects the fidelity and diversity of generated images. It measures the distance between the ground truth image distribution and the generated image distribution, where the distributions are constituted of Inception V3 [43] embeddings extracted from corresponding images. The CLIP score is used to evaluate the instruction-following ability of T2I models. It measures the similarity between the vision embeddings extracted from the generated image and text encoder embeddings extracted from corresponding captions.
@@ 4.1 Text-to-Image Generation
Text-to-image generation has become a cornerstone of modern artificial intelligence, enabling to create visual content based on textual descriptions. Pioneering models such as Generative Adversarial Networks (GANs) [10] make groundbreaking breakthroughs by generating high-fidelity images. AttnGAN [52] built on StackGAN [57] achieves better alignment with text instructions. However, GANs still faced challenges like training instability (e.g., mode collapse, where the model generates limited varieties of images) and difficulty with highly detailed or multi-object scenes, setting the stage for the next evolutionary step.
More recently, diffusion models [37, 13, 38] like Stable Diffusion [32] have emerged, creating realistic images by iteratively denoising random noise guided by text descriptions, setting a new standard for quality and versatility. However, the learning paradim and architectures diverge from well-established large language models (LLMs) [3],, making it difficult for artificial general intelligence featuring a shared framework for various modalities.
In this regard, a line of works [28, 29, 53] resort to autoregressive models for visual generation. Images are tokenized into a sequence of tokens and generated sequentially based on prefilled text tokens. Benefiting from large-scale models and datasets, they can create photorealistic images with a remarkable text-following capability. This field is further advanced by several open-source works, such as LlamaGen [42], Emu3 [48], and Janus [49].